OpenStudy (anonymous):

For which equation is x = 9 a solution? A. 3x – 7 = –34 B. x – 9 = 1 C. 2003-05-02-00-00_files/i0020000.jpg D. 4x + 5 = 41

OpenStudy (bee_see):

place x with 9 and see which one works.

OpenStudy (bee_see):

So for A, 3x-7, x=9. 3(9)-7=??? We should get -34. If not, then move on to B.

OpenStudy (anonymous):

Ok Thanks!

OpenStudy (bee_see):

what's the answer?

OpenStudy (anonymous):

To A it is -9 so I dont think that is right

OpenStudy (anonymous):

dosent work so on to c

OpenStudy (anonymous):

It is D

OpenStudy (anonymous):

I double checked

OpenStudy (bee_see):

right. Sorry A is 20. Yes, it's D.

OpenStudy (anonymous):

Its D 4*9+5=41

OpenStudy (anonymous):

Thanks!